The combined age of a mother and daughter is 32. in five years' time, the mother will be 5 times her daughter's age. How old are they now?
Answer by jorel1380(2518)

m+d=32
m+5=5(d+5)
m+5=5d+25
m=5d+20
5d+20=32-d
6d=12
d=2
m=30
The mother is 30. The daughter is 2..